# CounselBench: A Large-Scale Expert Evaluation and Adversarial Benchmark of Large Language Models in Mental Health Counseling

### v1

@@ 3.4 Expert Annotation Protocol
We implemented the evaluation rubric as a structured online questionnaire using Qualtrics, a wildly used platform for human subjects research. Prior to full deployment, we validated our questionnaire and study protocol with 8 participants through three rounds of pilot studies. Accordingly, we refined our evaluation questions and improved their clarity.
We recruited 100 licensed or professionally trained mental health practitioners across the U.S. through Upwork, the world’s largest talent platform commonly used in human-computer interaction (HCI) research for sourcing domain-expert annotators cscw-gig-economy-Upwork . All annotators met the requirement of having formal clinical counseling training or certification. Collectively, the annotators held 32 distinct types of licenses and degrees, and reported expertise spanning 43 specialized counseling areas; see Appendix B for more details. This broad coverage ensures that our evaluation reflects a diverse and representative cross-section of real-world counseling practice. Their credentials and clinical experience are summarized in Figure 2; additional demographics information is provided in Appendix B. Our annotator sample reflects the current demographics of the counseling profession in the U.S., which is predominantly composed of White and female practitioners McGuireMiranda2008 ; Verdieu2024 ; APA22 .
Each annotator was assigned a personalized survey containing 5 counseling questions, each paired with 4 responses: one written by an online human therapist and three generated by different LLMs. This resulted in 20 QA pairs per annotator. Each pair was independently evaluated by five different professionals, supporting inter-rater reliability and reducing individual bias. Importantly, for each question, all four responses were rated by the same group of annotators, enabling direct and fair comparison across models. In total, the study includes 2,000 annotated QA pairs.
Annotators were blind to the source of each response and were not informed that some had been generated by LLMs. Surveys were randomly sampled and individually assigned. Each annotator spent around 40 minutes per survey and was compensated \$20-\$35 depending on professional credentials and experience. This study was reviewed and approved as exempt research by the Institutional Review Board of University of Southern California (Study ID: UP-25-00022).
@@ 3.5 Expert Evaluation Results: Counseling Quality Across LLMs and Humans
We present the results of expert annotations comparing the counseling responses generated by GPT-4, LLaMA-3.3, Gemini-1.5-Pro, and online human therapists.
Figure 3: Average expert ratings.
Overall performance
Table 1 reports the average expert-annotated scores for each model, and Figure 3 provides a visual summary. Across most major evaluation dimensions, LLM-generated responses significantly outperformed those written by online human therapists. Among the models, LLaMA-3.3 received the highest overall ratings, leading on five of six dimensions. To assess significance, we performed Wilcoxon signed-rank tests wilcoxon_test_2005 , comparing the expert ratings of each LLM’s responses to those of online human therapists. We verified the consistency of our expert annotators’ ratings by calculating inter-rater reliability; annotator responses across measured metrics yielded Krippendorff’s alpha values greater than $0.7$, indicating a high level of inter-rater agreement inter-rater-reliability_2016 (See Table 2).
Table 1: Average expert ratings of counseling responses across six evaluation criteria. Each response was rated by five mental health professionals; scores were first averaged by question, then by model. Responses marked "I am not sure" were excluded. For Medical Advice, the percentage of “Yes” responses was computed per question (excluding “I’m not sure”) and averaged over all questions. Statistical significance was assessed using Wilcoxon signed-rank tests wilcoxon_test_2005 , comparing scores (averaged per question) between each LLM and online human therapists. Significance thresholds: $p^{*}<0.05$, $p^{**}<0.01$, $p^{***}<0.001$. Significance was not computed for Medical Advice, which was treated as a categorical variable.
Overall $\uparrow$ (range:1-5) Empathy $\uparrow$ (range: 1-5) Specificity $\uparrow$ (range: 1-5) Medical Advice (percentage of “Yes") Factual Consistency $\uparrow$ (range: 1-4) Toxicity $\downarrow$ (range: 1-5) GPT4     $3.28^{***}$     $3.37^{***}$     $3.46^{n.s.}$      0.07      $3.53^{***}$     $1.78^{***}$ Llama 3.3     $4.29^{***}$     $4.22^{***}$     $4.63^{***}$      0.14      $3.70^{***}$     $1.36^{***}$ Gemini-1.5-Pro     $3.26^{***}$     $2.76^{n.s.}$     $3.50^{n.s.}$      0.08      $3.52^{***}$     $1.64^{***}$ Online Human Therapists     2.60     2.72     3.29      0.17      2.92     2.56
Table 2: Averaged (over questions) Krippendorff’s alpha with ordinal weighting. Krippendorff’s alpha is not applied to binary Medical Advice.
@@
Despite strong scores, a higher proportion of LLaMA-3.3’s outputs (14%) were flagged for providing unauthorized medical advice (e.g., recommending therapy techniques), highlighting an important safety concern. In contrast, GPT-4 responses were more likely to include safety disclaimers, with approximately one-third of outputs explicitly declining to answer and recommending consultation with human professionals.
Qualitative Analysis of Annotator Rationales
To better understand expert judgments, we conducted a thematic analysis following established protocol Clarke_Braun_thematic_analysis_2017 . Two researchers on our team independently reviewed and thematically coded annotators’ open-text explanations. Seven primary issues among low-quality counseling responses are identified: 1) Lacking empathy or emotional attunement, 2) Displaying an inappropriate tone or attitude (e.g., dismissive, superficial), 3) Providing inaccurate suggestions (e.g., containing wrong information or making recommendations without sufficient evidence), 4) Offering unconstructive feedback (e.g., lacking clarity or actionability), 5) Demonstrating little personalization or relevance, 6) Containing language or terminology issues (e.g., typos, grammatical errors), and 7) Overgeneralizing or making judgments and assumptions without sufficient context.
We then prompted GPT-4.1 to label all low-rated responses (overall score $\leq$ 2) using these categories. Each response could exhibit multiple issues. A detailed breakdown is provided in Figure LABEL:fig:low_overall (Appendix LABEL:sub_appendeix_llm_judge_eval). The top two most frequently flagged failure types for each model are:
@@
Clinical Boundaries: Unauthorized Medical Advice in LLM responses
@@
To further examine safety risks, we analyzed the response excerpts flagged by expert annotators for containing medical advice. The most concerning identified issue was the inclusion of unauthorized clinical recommendations, especially prescribed psychotropic treatments (i.e., medications for mental health treatments, according to National Institute of Mental Health mental-health-medication-NIMH ). Among the LLMs, GPT-4 occasionally recommended specific medications, such as selective serotonin reuptake inhibitors (SSRIs). Annotators also identified that all three LLMs occasionally advised therapy techniques like cognitive behavioral therapy (CBT) or mindfulness. While these practices are common in clinical care, professionals emphasized that they require tailored application and oversight.
3.6 Can LLMs Reliably Judge Counseling Quality?
@@
Human expert evaluation provides gold-standard insight but is costly and difficult to scale gehrmann-etal-2021-gem . As LLMs are increasingly considered for deployment in mental health applications, one important question is whether models can serve as their own evaluators. Prior work has proposed the “LLM-as-Judge” paradigm for tasks such as summarization and factuality madaan2023selfrefine ; manakul2023selfcheckgpt , but its reliability in high-stakes, subjective domains like counseling remains unclear.
To investigate this, we tested eight advanced LLMs as automated judges, as listed in Figure 4. Each model was prompted with the same QA pairs used in Section 3.4, along with the evaluation criteria provided to human experts. Prompts followed the survey format used in our human annotation protocol, with minor formatting constraints tailored to each model; see Appendix LABEL:sub_appendeix_llm_judge_eval for details.
@@
Figure 4: Average evaluation scores across six dimensions (subplots) for counseling responses generated by GPT-4, LLaMA-3.3, Gemini-1.5-Pro, and online human therapists (x-axis in each subplot). Each colored line represents one evaluator, including eight LLM-based judges and human experts (red). Higher values indicate better performance except for Toxicity and Medical Advice. See Table LABEL:llm_judge_numeric_score for full numerical results.
@@
LLM Judges Overrate Responses and Miss Safety Failures. Figure 4 presents the scores provided by LLM judges, compared with human expert annotations. LLM-based evaluations consistently yielded inflated scores relative to human ratings. This was most pronounced for Factual Consistency, where LLM judges frequently assigned perfect or near-perfect scores to LLM-generated responses, regardless of their actual content. Ratings for Toxicity showed minimal variance across all sources. LLM judges uniformly assigned the lowest toxicity scores, even when human experts had flagged content as potentially harmful or inappropriate. This lack of sensitivity indicates a critical weakness in the use of LLMs for safety evaluations in high-risk domains.
LLM Rankings Diverge from Human Preferences.
@@
Beyond absolute scores, we compared model rankings based on average quality, as shown in Figure 4 and Table LABEL:llm_judge_numeric_score. The preferences expressed by LLM judges diverged sharply from those of human annotators. Most notably, Gemini-1.5-Pro was rated as the lowest-performing model by human expert annotators but was consistently ranked above GPT-4 by every LLM judge in terms of Overall Quality. This calls into question the reliability of LLMs in identifying relative quality, even under matched evaluation instructions.
LLM Judges Rarely Flag Problematic Text.
@@
We also tested whether LLM judges could identify problematic content at the sentence level. As in the human evaluation protocol, each model was asked to extract spans containing medical advice, factual errors, or toxic language. We compared the extracted text against sentences flagged by at least two human expert annotators, treating these as widely agreed indicators of failure.
To standardize comparison, we mapped all annotations to complete sentences; see Appendix LABEL:sec:sentence_level for details. Table 3 reports the proportion of widely-flagged sentences detected by each LLM judge. Most LLMs failed to detect even a single flagged toxic or incorrect sentence. Some models detected limited instances of unauthorized medical advice, but none showed robust or reliable detection across safety categories. These results help explain the inflated consistency and toxicity scores in earlier evaluations and highlight the limitations of LLM-based judging in high-stakes settings.
Table 3: Percentage of widely flagged sentences (labeled by at least two annotators) detected by each LLM judge. Rows correspond to LLM judges, and columns correspond to four response sources: GPT-4, LLaMA-3.3, Gemini-1.5-Pro, and online human therapists. Higher values reflect closer alignment with expert annotations.
Flagged Medical Advice Flagged Factual Errors Flagged Toxic Content Judges GPT-4 Llama-3 Gemini Human GPT-4 Llama-3 Gemini Human GPT-4 Llama-3 Gemini Human GPT-3.5-turbo 0 0 0 0 0 0 9.1% 8% 0 0 10% 2.4% GPT-4 25% 11.1% 11.1% 23.1% 0 0 0 5.3% 0 0 0 0 Llama-3.1 37.5% 5.6% 22.2% 30.8% 0 0 0 8% 0 0 0 0 Llama-3.3 37.5% 11.1% 33.3% 34.6% 0 0 9.1% 10.7% 0 0 0 0 Claude-3.5 25% 11.1% 11.1% 3.9% 0 0 0 6.7% 0 0 0 0 Claude-3.7 37.5% 11.1% 33.3% 7.7% 0 0 9.1% 8% 0 0 0 0 Gemini-1.5-Pro 12.5% 5.6% 33.3% 15.4% 0 0 0 61.3% 0 0 0 0 Gemini-2.0-Flash 37.5% 11.1% 11.1% 26.9% 0 0 0 6.7% 0 0 0 0
4 CounselBench-ADV: An Adversarial Benchmark for Surfacing LLM Failures in Counseling
@@
While expert annotators rated LLM-generated counseling responses as generally high quality, our in-depth analysis revealed recurring issues, such as inappropriate advice or generic feedback (§3.5). To probe these failure modes more systematically, we introduce CounselBench-ADV, an adversarial benchmark constructed by mental health professionals.
Identifying Fine-grained Failure Modes
@@
To construct a targeted adversarial benchmark, we first need to isolate concrete failure modes that LLMs exhibit in realistic counseling settings. Broad categories (e.g., "lacking empathy" and "lacking personalization") identified in Section 3.5 are too general to support targeted adversarial construction. To enable precise probing, we conducted a deeper review of expert rationales and flagged responses to extract six concrete failure modes:
  • GPT-4: The response provides specific medication (1. medication) and suggests specific therapy techniques suggestions (2. therapy).
  • Llama-3.3: The response speculates about medical symptoms (3. symptoms) and is judgmental (4. judmental).
  • Gemini-1.5-Pro: The response is apathetic (5. apathetic) and is based on unsupported assumptions (6. assumptions).
@@
Here we selected two per model that best represent the kinds of specific issues each model; see Table LABEL:tab:six_issues for details. Each issue is linked to real examples from CounselBench-EVAL, providing clear foundations for clinicians to construct new questions that elicit these failures in a controlled and measurable way.
Adversarial Dataset Construction
@@
We rehired 10 mental health professionals who previously participated as expert annotators in CounselBench-EVAL. Each clinician was provided with detailed definitions and real examples illustrating each failure type, drawn directly from flagged responses in CounselBench-EVAL (Table LABEL:tab:six_issues). They were instructed to independently author two realistic, single-turn counseling questions for each failure mode, i.e., crafted in such a way that it would be plausible for an LLM to fall into the corresponding error.
Each clinician authored 12 adversarial questions (2 per issue × 6 issues), resulting in a total of 120 adversarial questions. Clinicians were asked to ensure the realism and diversity of the questions, reflecting what actual patient might ask during counseling. Importantly, these questions do not directly contain the failure, but are designed to trigger the failure in LLM responses. The resulting dataset provides a high-precision probe of specific vulnerabilities in LLM-generated counseling.
Eliciting and Evaluating Model Failures
To evaluate model behavior under adversarial conditions, we used the same prompting setup as in CounselBench-EVAL3.2) to generate counseling responses for each of the 120 adversarial questions. Each prompt was submitted to eight LLMs in Table 4. For each model, we issued the same prompt three times to sample variations in generation. In total, this yielded 2,880 QA pairs (120 questions × 8 models × 3 generations).
We used a hybrid annotation protocol to determine whether each adversarial question successfully triggered its target failure mode. One mental health professional reviewed a random subset of 96 QA pairs (3% of the full set), while GPT-4.1 was used to annotate the full set. We provided GPT-4.1 with detailed definitions of each failure type and included an in-context example drawn from CounselBench-EVALLABEL:sec:aversarial_by_gpt_41). In this way, the GPT evaluator had both conceptual clarity and concrete guidance to support consistent and accurate labeling. Agreement between the expert and GPT annotator reached 72.9%, indicating high reliability inter-rater-reliability_2016 .
@@
Table 4: Fraction of responses identified to contain each targeted failure mode. Higher values reflect greater model vulnerability to the targeted issue.
@@
Specific Issue GPT-3.5-Turbo GPT-4 Llama-3.1 Llama-3.3 Claude-3.5-Sonnet Claude-3.7-Sonnet Gemini-1.5-Pro Gemini-2.0-Flash 1. Medication   0.25   0.32   0.40   0.55   0.47   0.53   0.42   0.47 2. Therapy   0.28   0.25   0.60   0.53   0.50   0.53   0.30   0.35 3. Symptoms   0.20   0.82   0.83   0.83   0.87   0.83   0.67   0.68 4. Judgmental   0.12   0.02   0.22   0.22   0.15   0.08   0.15   0.08 5. Apathetic   0.85   0.22   0.18   0.13   0.10   0.12   0.47   0.42 6. Assumptions   0.53   0.35   0.73   0.80   0.57   0.32   0.42   0.48
Adversarial Questions Effectively Trigger Targeted Failures
@@
Table 4 reports the frequency of each failure mode across model outputs based on GPT-4.1 evaluation. The adversarial questions effectively elicited the intended issues in a large fraction of cases. Notably, speculation about medical symptoms emerged as the most prevalent failure: it was triggered in 82% of GPT-4 responses, 83% of LLaMA series responses, over 83% of Claude models, and more than 67% of Gemini outputs. This pattern raises serious concerns, as such speculation can lead to inappropriate clinical guidance without proper diagnosis. In contrast, judgmental tone was the least common issues across models, appearing in only 2% of GPT-4 responses and 22% of responses from the LLaMA series.

### v2

@@ 4.4 Expert Annotation Protocol
We implemented the evaluation rubric as a structured online questionnaire using Qualtrics, a widely used platform for human subjects research.
Pilot Studies. Prior to full deployment, we validated our questionnaire and study protocol with 8 participants through three rounds of pilot studies. Accordingly, we refined our evaluation questions and improved their clarity. For example, some experts occasionally expressed uncertainty on Medical Advice and Factual Consistency, which is not observed in other dimensions, so we included the option ”I am not sure” in the main study to reduce noise and avoid forcing uncertain judgments.
Recruitment and Credential Verifications. We recruited 100 licensed or professionally trained mental health practitioners across the U.S. through Upwork, the world’s largest talent platform commonly used in human-computer interaction (HCI) research for sourcing domain-expert annotators (Fulker & Riedl, 2024). All annotators had formal counseling training or certification, and we individually verified their education, credentials, or professional experience (Appendix B).
Annotation Protocol. Each annotator was assigned a randomly sampled survey containing 5 counseling questions, each paired with 4 responses: one from an online human therapist and three from different LLMs. The four responses for each question are in randomized order to mitigate position bias. This resulted in 20 QA pairs per annotator. Each pair was independently evaluated by five different professionals, supporting inter-rater reliability and reducing individual bias. For each question, all four responses were rated by the same group of annotators, enabling direct and fair comparison across models. In total, the study includes 2,000 annotated QA pairs. Annotators were blind to the source of each response and were not informed that some had been generated by LLMs.
Annotator Demographics and Engagement Collectively, the 100 annotators held 32 distinct types of licenses and degrees, and reported expertise spanning 43 specialized counseling areas (summarized in Figure 2; more details in Appendix B). This broad coverage ensures that our evaluation reflects a diverse and representative cross-section of real-world practice. The demographic composition of our annotator sample aligns with national statistics for the U.S. counseling profession, which shows a higher proportion of white and female practitioners (McGuire & Miranda, 2008; Verdieu, 2024; American Psychological Association, 2022). The median written rationale length was 576.5 words across 20 QA pairs (IQR: 339.8 - 866.5), and the median time spent was 1 hr 22 min (IQR: 52 min - 3 hrs 9 min), indicating substantial engagement.
@@ 4.5 Expert Evaluation Results: Counseling Quality Across LLMs and Humans
We present the results of expert annotations comparing the responses generated by GPT-4, LLaMA-3.3, Gemini-1.5-Pro, and online human therapists.
Overall Performance. Table 1 reports the average expert-annotated scores for each model. Among the models, LLaMA-3.3 received the highest overall ratings, leading on five of six dimensions. Pairwise significance tests (Woolson, 2005) are reported in Appendix LABEL:appendix_pairwise_significance_test. Despite strong scores, a higher proportion of LLaMA-3.3’s outputs (14%) were flagged for providing unauthorized medical advice (e.g., recommending therapy techniques), highlighting an important safety concern. In contrast, GPT-4 responses were more likely to include safety disclaimers, with approximately one-third of outputs explicitly declining to answer and recommending consultation with human professionals.
Table 1: Average expert ratings of counseling responses across six evaluation criteria. Each response was rated by five mental health professionals; scores were first averaged by question, then by model. Responses marked ”I am not sure” were excluded. For Medical Advice, the percentage of “Yes” responses was computed per question (excluding “I’m not sure”) and averaged over all questions.
Overall $\uparrow$ (range:1-5) Empathy $\uparrow$ (range: 1-5) Specificity $\uparrow$ (range: 1-5) Medical Advice (percentage of “Yes”) Factual Consistency $\uparrow$ (range: 1-4) Toxicity $\downarrow$ (range: 1-5) GPT4     $3.28$     $3.37$     $3.46$      0.07      $3.53$     $1.78$ Llama 3.3     $4.29$     $4.22$     $4.63$      0.14      $3.70$     $1.36$ Gemini-1.5-Pro     $3.26$     $2.76$     $3.50$      0.08      $3.52$     $1.64$ Online Human Therapists     2.60     2.72     3.29      0.17      2.92     2.56
Table 2: Mean (over questions) Krippendorff’s alpha (ordinal); not applied to binary Medical Advice.
@@
Annotation Consistency. We verified the consistency of our expert annotators’ ratings by calculating inter-rater reliability (Table 2); annotator responses across measured metrics yielded Krippendorff’s alpha values  $\geq 0.7$, indicating a high level of inter-rater agreement (Zapf et al., 2016).
Annotation Analysis. Additional analyses of annotator behavior and score distributions are reported in Appendix C. We find that annotators who spent more time (top 50%) wrote longer rationales and labeled more sentences as Medical Advice. Significant differences in Empathy, Medical Advice, and Factual Consistency were also observed across annotators with different years of experience. Finally, Table LABEL:tab:uncertain-percentage reports the fraction of “I am not sure” selections, which were infrequent overall.
Qualitative Analysis of Annotator Rationales To better understand expert judgments, we conducted a thematic analysis following established protocol (Clarke & Braun, 2017). Two researchers on our team independently reviewed and thematically coded annotators’ open-text explanations. Seven primary issues among low-quality responses are identified (listed in Appendix LABEL:sub_appendeix_llm_judge_eval). We then prompted GPT-4.1 to label all low-rated responses (overall score $\leq$ 2) using these categories. Each response could exhibit multiple issues. A detailed breakdown is provided in Figure LABEL:fig:low_overall (Appendix LABEL:sub_appendeix_llm_judge_eval). The top two most frequently flagged failure types for each model are:
@@
Clinical Boundaries: Unauthorized Medical Advice in LLM responses To further examine safety risks, we analyzed the response excerpts flagged by expert annotators for containing medical advice. The most concerning identified issue was the inclusion of unauthorized clinical recommendations, especially prescribed psychotropic treatments (i.e., medications for mental health treatments, according to National Institute of Mental Health (of Mental Health, 2025)).
Among the LLMs, GPT-4 occasionally recommended specific medications, such as selective serotonin reuptake inhibitors (SSRIs). Annotators also identified that all three LLMs occasionally advised therapy techniques like cognitive behavioral therapy (CBT) or mindfulness. While these practices are common in clinical care, professionals emphasized that they require tailored application and oversight. In CounselBench-ADV, we further refined this category into (i) recommending specific medications, (ii) suggesting therapy techniques, and (iii) speculating about medical symptoms.
4.6 Can LLMs Reliably Judge the Quality of Responses?
@@
Human expert evaluation provides gold-standard insight but is costly and difficult to scale (Gehrmann et al., 2021). As LLMs are increasingly considered for deployment in mental health applications, one important question is whether models can serve as their own evaluators. Prior work has proposed the “LLM-as-Judge” paradigm for tasks such as summarization and factuality (Madaan et al., 2023; Manakul et al., 2023), but its reliability in high-stakes, subjective domains like mental health remains unclear. To investigate this, we tested eight advanced LLMs as automated judges, as listed in Figure 3. Each model was prompted with the same QA pairs used in Section 4.4, along with the evaluation criteria provided to human experts. Prompts followed the survey format used in our human annotation protocol, with minor formatting constraints tailored to each model (see Appendix LABEL:sub_appendeix_llm_judge_eval), and additional results for LLM judges are in Appendix LABEL:llm_judge_additional.
@@
Figure 3: Average evaluation scores across six dimensions (subplots) for counseling responses generated by GPT-4, LLaMA-3.3, Gemini-1.5-Pro, and online human therapists (x-axis in each subplot). Each colored line represents one evaluator, including nine LLM-based judges and human experts (red). Higher values indicate better performance except for Toxicity and Medical Advice. See Table LABEL:llm_judge_numeric_score for full numerical results.
@@
LLM Judges Overrate Responses and Miss Safety Failures. Figure 3 presents the scores provided by LLM judges, compared with human expert annotations. Most LLM-based evaluations yielded inflated scores relative to human ratings. This was most pronounced for Factual Consistency, where LLM judges frequently assigned perfect or near-perfect scores to LLM-generated responses, regardless of their actual content. Ratings for Toxicity showed minimal variance across all sources. All LLM judges uniformly assigned the lowest toxicity scores, even when human experts had flagged content as potentially harmful or inappropriate. This lack of sensitivity indicates a critical weakness in the use of LLMs for safety evaluations in high-risk domains.
@@
LLM Rankings Diverge from Human Preferences. Beyond absolute scores, we compared model rankings based on average quality, as shown in Figure 3 and Figure 3. The preferences expressed by LLM judges diverged sharply from those of human annotators, except for GPT-5 (dark blue line in Table 3), which calls into question the reliability of LLMs in identifying relative quality. Most notably, Gemini-1.5-Pro was rated as the lowest-performing model by human expert annotators but was consistently ranked above GPT-4 by every LLM judge for Overall Quality.
@@
LLM Judges Rarely Flag Problematic Text. We also tested whether LLM judges could identify problematic content at the sentence level. As in the human evaluation protocol, each model was asked to extract spans containing medical advice, factual errors, or toxic language. We compared the extracted text against sentences flagged by at least two human annotators, treating these as widely agreed indicators of failure (more details in Appendix LABEL:sec:sentence_level). Table LABEL:tab:miss_percentage in the Appendix reports the proportion of widely-flagged sentences detected by each LLM judge. Most models failed to catch any toxic or incorrect sentences. Some caught limited instances of unauthorized medical advice, but none showed robust or reliable detection across safety categories. These results help explain the inflated consistency and toxicity scores and highlight the limitations of LLM-based judging.
5 CounselBench-ADV: An Adversarial Benchmark for Surfacing LLM Failures
@@
While expert annotators rated LLM-generated responses to open-ended mental health questions as generally high quality, our in-depth analysis revealed recurring issues, such as inappropriate advice or generic feedback (§4.5). To probe these failure modes more systematically, we introduce CounselBench-ADV, an adversarial benchmark constructed by mental health professionals.
@@
Identifying Fine-grained Failure Modes To construct a targeted adversarial benchmark, we first need to isolate that LLMs exhibit. Broad categories (e.g., “lacking personalization”) from Section 4.5 are too general. To enable precise probing, we conducted a deeper review of expert rationales and flagged responses to extract concrete failure modes:
  • GPT-4: provides specific medication (1. medication) and suggests specific therapy techniques (2. therapy).
  • Llama-3.3: speculates about medical symptoms (3. symptoms) and is judgmental (4. judmental).
  • Gemini-1.5-Pro: is apathetic (5. apathetic) and is based on unsupported assumptions (6. assumptions).
@@
Here we selected two per model that best represent the kinds of specific issues each model(see Table LABEL:tab:six_issues in Appendix LABEL:appendix_failure_modes). Each issue is linked to real examples from CounselBench-EVAL, providing clear foundations for clinicians to construct new questions in a controlled and measurable way.
@@
Adversarial Dataset Construction We rehired 10 mental health professionals who previously participated in CounselBench-EVAL. Each clinician was provided with detailed definitions and real examples illustrating each failure type, drawn directly from flagged responses in CounselBench-EVAL (Table LABEL:tab:six_issues). They were instructed to independently author two realistic mental health questions for each failure mode, i.e., crafted in such a way that it would be plausible for an LLM to fall into the corresponding error. Each clinician authored 12 adversarial questions (2 per issue × 6 issues), resulting in a total of 120 questions. Clinicians were asked to ensure the realism and diversity of the questions, reflecting what actual patient might ask. Importantly, these questions do not directly contain the failure, but are designed to trigger the failure in LLM responses. The resulting dataset provides a high-precision probe of specific vulnerabilities in LLM-generated responses.
Evaluation Setup. We used the same prompting setup as in CounselBench-EVAL for these adversarial questions with nine LLMs (results are in Table 3). For each model, we issued the same prompt three times to sample variations in generation. In total, this yielded 3,240 QA pairs (120 questions × 9 models × 3 generations). To determine whether a response triggered its targeted failure mode, we used GPT-4.1 as a scalable labeling tool. The task was deliberately framed as binary classification with clear clinician-authored definitions and one in-context example per failure type (§LABEL:sec:aversarial_by_gpt_41), making it simpler and more reproducible than multi-dimensional quality scoring.
Human Validation. A mental health professional spot-checked 96 responses, yielding 72.9% agreement with GPT-4.1 (Stemler, 2004). This level of concordance indicates that GPT-4.1 can serve as a practical proxy for detecting whether a predefined failure was triggered, though it is not a substitute for full expert evaluation. We therefore use GPT-based labels only to quantify trigger rates in CounselBench-Adv, while all primary analyses of quality and safety remain grounded in expert annotations from CounselBench-Eval. Expanding human annotation to cover the full set of 3,240 adversarial responses represents an important direction for future work.
@@
Table 3: Fraction of responses identified to contain each targeted failure mode. Higher values reflect greater model vulnerability to the targeted issue. Red: zero-shot; Orange: few-shot.
@@
Specific Issue GPT-3.5-Turbo GPT-4 GPT-5 Llama-3.1 Llama-3.3 Claude-3.5-Sonnet Claude-3.7-Sonnet Gemini-1.5-Pro Gemini-2.0-Flash 1. Medication   0.25   0.32   0.97   0.40   0.55   0.47   0.53   0.42   0.47 2. Therapy   0.28   0.25   0.93   0.60   0.53   0.50   0.53   0.30   0.35 3. Symptoms   0.20   0.82   0.85   0.83   0.83   0.87   0.83   0.67   0.68 4. Judgmental   0.12   0.02   0.02   0.22   0.22   0.15   0.08   0.15   0.08 5. Apathetic   0.85   0.22 0   0.18   0.13   0.10   0.12   0.47   0.42 6. Assumptions   0.53   0.35   0.17   0.73   0.80   0.57   0.32   0.42   0.48 1. Medication   0.25   0.47   0.92   0.45   0.48   0.62   0.62   0.50   0.57 2. Therapy   0.43   0.40   0.95   0.78   0.75   0.65   0.58   0.40   0.50 3. Symptoms   0.42   0.83   0.85   0.82   0.82   0.85   0.82   0.72   0.65 4. Judgmental   0.08   0.08   0.07   0.22   0.20   0.15   0.17   0.05   0.15 5. Apathetic   0.42   0.07 0   0.17   0.07   0.08   0.03   0.12   0.50 6. Assumptions   0.45   0.67   0.15   0.80   0.82   0.67   0.55   0.42   0.65
@@
Adversarial Questions Effectively Trigger Targeted Failures Table 3 reports the frequency of each failure mode across model outputs based on GPT-4.1 evaluation. The adversarial questions effectively elicited the intended issues in a large fraction of cases. Notably, speculation about medical symptoms emerged as the most prevalent failure: it was triggered in 82% of GPT-4 responses, 83% of LLaMA series responses, over 83% of Claude models, and more than 67% of Gemini outputs. This pattern raises serious concerns, as such speculation can lead to inappropriate clinical guidance without proper diagnosis. In contrast, judgmental or apathetic tone was the less common issues across models.
To further explore potential causes and mitigation strategies, we conducted a follow-up experiment on CounselBench-ADV using few-shot prompting (details described in Appendix LABEL:sec:few-shot-adv). The resulting failure rates are also shown in Table 3. Across most categories, few-shot prompting yielded limited improvements. This suggests that many failure modes are likely rooted in pretraining and not easily resolved through in-context examples.